OpenStudy (anonymous):

Help? Solve the triangle. B = 73°, b = 15, c = 8

OpenStudy (anonymous):

 (Assuming it's not a right triangle) We can use the law of sines to solve for C. \[\frac{\sin B}{b} = \frac{\sin C}{c}\] Plug in your values and solve for C. (Remember to use your calculator in degrees mode.)

OpenStudy (anonymous):

What the heck? Why does my school say to do b/sinB=c/sinC? Or does it do the same thing in the end?

OpenStudy (anonymous):

They do the same thing in the end. It doesn't matter, as long as it's consistent. Keep the sines on the same side of the fraction.

OpenStudy (anonymous):